OpenStudy (anonymous):

Help please i will fan Which of the following is equal to square root of the cube root of 2? 2 to the power of 2 over 3 2 to the power of 3 over 2 2 to the power of 1 over 3 2 to the power of 1 over 6

OpenStudy (anonymous):

@MeganChase

OpenStudy (anonymous):

give me a minute

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i really dont know i can help sorry

OpenStudy (anonymous):

(2^1/3)^1/2 = 2^1/6 (a^b)^c = a^(bc)

OpenStudy (anonymous):

\[\sqrt{2^{\frac{ 1 }{ 3 }}}=\left( 2^{\frac{ 1 }{ 3}} \right)^{\frac{ 1 }{ 2 }}=2^{\frac{ 1 }{ 3 }\times \frac{ 1 }{ 2 }}=2^{\frac{ 1 }{ 6 }}\]

OpenStudy (anonymous):

thanks guys

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

give me a sec and i fan you both